Help me solve by defining a variable, writing an equation, and solving
A 10-foot ladder is set 6 feet from the base of the wall. How far up the wall will the ladder reach?

Answer by nerdybill(7384)

A 10-foot ladder is set 6 feet from the base of the wall. How far up the wall will the ladder reach?
.
The wall, ground and the ladder forms a "right triangle" (where the hypotenuse is the ladder).
.
Let x = distance the ladder will reach on the wall
then
x^2 + 6^2 = 10^2
x^2 + 36 = 100
x^2 = 64
x = 8 feet